# Why Does Little Robustness Help? A Further Step Towards Understanding Adversarial Transferability

### v7

@@ 6 Analyses and Discussions
To validate this standpoint, we show that with a different training method such as SAM, the transferability of ViT can be significantly improved. Our experiments are conducted on well-trained ImageNet classifiers, all publicly available from PyTorch image model library timm [71]. We consider standard ViT surrogates ViT-B/16 and ViT-B/32, which are both first pre-trained on a large-scale dataset and then fine-tuned on ImageNet using a combination of MU and RandAugment [8]. In contrast, their SAM versions ViT-B/16-SAM and ViT-B/32-SAM are merely trained from scratch. We use both normally and adversarially trained CNN classifiers as target models. We craft AEs on these 4 surrogates using both standard PGD and DIM under $\epsilon=16/255$ on the ImageNet-compatible dataset as in [49]. The accuracy of these surrogates (referring to [71]) and ASRs are reported in Tab. IX. The results indicate that standard ViT models have much weaker transferability than their SAM versions, despite the fact that they have higher accuracy.
Q6: Will the transferability benefit of better surrogate in “lab” generalize to “real-world” environment? Recently, Mao et al. [42] suggests that many conclusions about adversarial transferability done in the simplistic controlled “lab” environments may not hold in the real world. We thus conduct experiments to verify the generality of our conclusions using AEs crafted against our ImageNette surrogates.
Specifically, aligning with Mao et al. [42], we conduct transfer attacks on three commercial MLaaS platforms—AWS Rekognition, Alibaba Cloud, and Baidu Cloud, using their code, and report the results in Tab. X. The results are generally consistent with that in Tab. III. Particularly, we highlight here that our experimental results on MLaaS platforms contradict the observation in Mao et al. [42] that “adversarial training and data augmentation do not show strong correlations to the transfer attack”. Tab. X indicates that strong data augmentations like CM and MU generally exhibit lower ASRs than ST for both targeted and untargeted attacks. In addition, we can observe that “little robustness” (see AT, $\epsilon=1$) improves transferability while large adversarial magnitude (see AT, $\epsilon=5$) degrades it.
Q7: How does SAM yield superior alignment towards every training solution? Given its superior performance on transferability and gradient alignment, it is natural to wonder why SAM exhibits this property. Recent works [22, 33] empirically demonstrate that attacking an ensemble of models in the distribution obtained via Bayesian learning substantially improves the transferability.
@@
Another concurrent work [44] theoretically establishes SAM as an optimal relaxation of Bayes object. Consequently, this suggests that the transferability benefit of SAM may owe to the fact it optimizes a single model that represents an ensemble (expectation) of models, i.e., attacking a SAM solution is somewhat equivalent to attacking an ensemble of solutions. We conjecture that the strong input gradient alignment towards every training solution may reflect this “expectation” effect of SAM (Bayes). Notably, our interpretation here is merely intuitive and speculative, however, SAM still remains poorly understood to date. We believe that uncovering the exact reason is far beyond the scope of our study, while our experimental observations provide valuable insights and may contribute to a better understanding of SAM for future study.
@@ 7.1 Beyond Adversarial Transferability
Typically, Liang et al. [35] have shed light on the bidirectional relationship between these two forms of transferability. They presented theoretical and empirical results showing that domain transferability implies adversarial transferability, and vice versa. This finding suggests that approaches aimed at enhancing domain transferability during model training may also have a positive impact on adversarial transferability.
Within the line of research [81, 69, 83, 65] investigating factors influencing domain transferability, several works [79, 53, 63] have focused on the impact of adversarial training. These works consistently showed that adversarially trained models transfer better than non-robust models on tasks in different target domains, i.e., adversarial training improves domain transferability, thus asserting a connection between adversarial robustness and domain transferability.
However, recently, Xu et al. [77] challenged this connection and contended that robustness alone is neither necessary nor sufficient for domain transferability. Through comprehensive experiments, they demonstrated that models with better robustness do not necessarily guarantee enhanced domain transferability. Instead, the regularization effect of adversarial training offers a fundamental explanation for the relationship between adversarial training and domain transferability, namely, adversarial training improves domain transferability through its regularization effect.
Taken together, the conclusions of these researches suggest that adversarial training is expected to improve both two kinds of transferability with its regularization effect. However, in this work, though we find the regularization effect induced by adversarial training does play a crucial role in improving model smoothness, leading to enhanced adversarial transferability, the situation is much more complex: adversarially trained models with small budgets yield better adversarial transferability, large budgets perform worse [57, 58], and yet, these same robust models does uniformly transfer better in downstream tasks [53].
Our work dissects this nuanced relationship and establishes a connection between the drawbacks of adversarial training in adversarial transferability and the data distribution shift. From the distribution perspective, the key distinction between these two kinds of transferability is that domain transferability does not necessarily require the model to match the source domain perfectly since the target domain is independent of it, yet successful untargeted (targeted) adversarial attacks owe to pulling (pushing) the generated AEs out of (into) the original (target) sample distribution [89], thus demanding the models fitting the data domain well.
@@ 7.2 Beyond Surrogate Training
Intriguingly, besides the perspective of surrogate training, we notice that some concurrent works adopt similar design philosophies as ours from other perspectives to improve adversarial transferability. Wu et al. [74] and Ge et al. [19] both suggest a regularization term that encourages small loss gradient norms in the input space during AE generation, while the latter also performs gradient averaging on multiple randomly sampled neighbors. On the other hand, Chen et al. [5] study the model ensemble perspective and propose a common weakness attack, which adopts SAM and a cosine similarity encourager to optimize AEs, aiming to converge to points close to the flat local optimum of each model. These works also demonstrate the effectiveness of optimizing smoothness and similarity for improving adversarial transferability.
Other approaches that study from the surrogate perspective transform the model at the test time (i.e., after training, when performing the attack). Ghost network [34] attacks a set of ghost networks generated by densely applying dropout at intermediate features. SGM [72] incorporates more gradients from the skip connections of ResNet to generate AEs, while LinBP [24] and BPA [66] both restore the truncated gradient caused by non-linear layers.
Our work complements them all since we view adversarial transferability as an inherent property of the surrogate model itself and propose methods to construct surrogate models specialized for transfer attacks at the training time. We leave the exploration and discussion regarding a unified framework [67] that can reason adversarial transferability from all perspectives for future work.

### v8

@@ 6 Analyses and Discussions
To validate this standpoint, we show that with a different training method such as SAM, the transferability of ViT can be significantly improved. Our experiments are conducted on well-trained ImageNet classifiers, all publicly available from PyTorch image model library timm [rw2019timm]. We consider standard ViT surrogates ViT-B/16 and ViT-B/32, which are both first pre-trained on a large-scale dataset and then fine-tuned on ImageNet using a combination of MU and RandAugment [Cubuk_2020_CVPR_Workshops]. In contrast, their SAM versions ViT-B/16-SAM and ViT-B/32-SAM are merely trained from scratch. We use both normally and adversarially trained CNN classifiers as target models. We craft AEs on these 4 surrogates using both standard PGD and DIM under $\epsilon=16/255$ on the ImageNet-compatible dataset as in [qin2022boosting]. The accuracy of these surrogates (referring to [rw2019timm]) and ASRs are reported in Tab. IX. The results indicate that standard ViT models have much weaker transferability than their SAM versions, despite the fact that they have higher accuracy.
Q6: Will the transferability benefit of better surrogate in “lab” generalize to “real-world” environment? Recently, Mao et al. [maoP2022TransferAttacks2022] suggests that many conclusions about adversarial transferability done in the simplistic controlled “lab” environments may not hold in the real world. We thus conduct experiments to verify the generality of our conclusions using AEs crafted against our ImageNette surrogates.
Specifically, aligning with Mao et al. [maoP2022TransferAttacks2022], we conduct transfer attacks on three commercial MLaaS platforms—AWS Rekognition, Alibaba Cloud, and Baidu Cloud, using their code, and report the results in Tab. X. The results are generally consistent with that in Tab. III. Particularly, we highlight here that our experimental results on MLaaS platforms contradict the observation in Mao et al. [maoP2022TransferAttacks2022] that “adversarial training and data augmentation do not show strong correlations to the transfer attack”. Tab. X indicates that strong data augmentations like CM and MU generally exhibit lower ASRs than ST for both targeted and untargeted attacks. In addition, we can observe that “little robustness” (see AT, $\epsilon=1$) improves transferability while large adversarial magnitude (see AT, $\epsilon=5$) degrades it.
Q7: How does SAM yield superior alignment towards every training solution? Given its superior performance on transferability and gradient alignment, it is natural to wonder why SAM exhibits this property. Recent works [pmlr-v180-gubri22a, li2023making] empirically demonstrate that attacking an ensemble of models in the distribution obtained via Bayesian learning substantially improves the transferability.
@@
Another concurrent work [mollenhoff2023sam] theoretically establishes SAM as an optimal relaxation of Bayes object. Consequently, this suggests that the transferability benefit of SAM may owe to the fact it optimizes a single model that represents an ensemble (expectation) of models, i.e., attacking a SAM solution is somewhat equivalent to attacking an ensemble of solutions. We conjecture that the strong input gradient alignment towards every training solution may reflect this “expectation” effect of SAM (Bayes). Notably, our interpretation here is merely intuitive and speculative, however, SAM still remains poorly understood to date. We believe that uncovering the exact reason is far beyond the scope of our study, while our experimental observations provide valuable insights and may contribute to a better understanding of SAM for future study.
@@ 7.1 Beyond Adversarial Transferability
Typically, Liang et al. [liang2021uncovering] have shed light on the bidirectional relationship between these two forms of transferability. They presented theoretical and empirical results showing that domain transferability implies adversarial transferability, and vice versa. This finding suggests that approaches aimed at enhancing domain transferability during model training may also have a positive impact on adversarial transferability.
Within the line of research [yosinski2014transferable, wang2019characterizing, zhang2021quantifying, wang2022generalizing] investigating factors influencing domain transferability, several works [yi2021improved, salman2020adversarially, utrera2021adversariallytrained] have focused on the impact of adversarial training. These works consistently showed that adversarially trained models transfer better than non-robust models on tasks in different target domains, i.e., adversarial training improves domain transferability, thus asserting a connection between adversarial robustness and domain transferability.
However, recently, Xu et al. [xuICML2022Adversarially2022] challenged this connection and contended that robustness alone is neither necessary nor sufficient for domain transferability. Through comprehensive experiments, they demonstrated that models with better robustness do not necessarily guarantee enhanced domain transferability. Instead, the regularization effect of adversarial training offers a fundamental explanation for the relationship between adversarial training and domain transferability, namely, adversarial training improves domain transferability through its regularization effect.
Taken together, the conclusions of these researches suggest that adversarial training is expected to improve both two kinds of transferability with its regularization effect. However, in this work, though we find the regularization effect induced by adversarial training does play a crucial role in improving model smoothness, leading to enhanced adversarial transferability, the situation is much more complex: adversarially trained models with small budgets yield better adversarial transferability, large budgets perform worse [springer2021little, springer2021uncovering], and yet, these same robust models does uniformly transfer better in downstream tasks [salman2020adversarially].
Our work dissects this nuanced relationship and establishes a connection between the drawbacks of adversarial training in adversarial transferability and the data distribution shift. From the distribution perspective, the key distinction between these two kinds of transferability is that domain transferability does not necessarily require the model to match the source domain perfectly since the target domain is independent of it, yet successful untargeted (targeted) adversarial attacks owe to pulling (pushing) the generated AEs out of (into) the original (target) sample distribution [zhu2022toward], thus demanding the models fitting the data domain well.
@@ 7.2 Beyond Surrogate Training
Intriguingly, besides the perspective of surrogate training, we notice that some concurrent works adopt similar design philosophies as ours from other perspectives to improve adversarial transferability. Wu et al. [wu2023gnp] and Ge et al. [ge2023boosting] both suggest a regularization term that encourages small loss gradient norms in the input space during AE generation, while the latter also performs gradient averaging on multiple randomly sampled neighbors. On the other hand, Chen et al. [chen2023rethinking] study the model ensemble perspective and propose a common weakness attack, which adopts SAM and a cosine similarity encourager to optimize AEs, aiming to converge to points close to the flat local optimum of each model. These works also demonstrate the effectiveness of optimizing smoothness and similarity for improving adversarial transferability.
Other approaches that study from the surrogate perspective transform the model at the test time (i.e., after training, when performing the attack). Ghost network [li2020learning] attacks a set of ghost networks generated by densely applying dropout at intermediate features. SGM [Wu2020Skip] incorporates more gradients from the skip connections of ResNet to generate AEs, while LinBP [NEURIPS2020_00e26af6] and BPA [wang2023rethinking] both restore the truncated gradient caused by non-linear layers.
Our work complements them all since we view adversarial transferability as an inherent property of the surrogate model itself and propose methods to construct surrogate models specialized for transfer attacks at the training time. We leave the exploration and discussion regarding a unified framework [wang2021a] that can reason adversarial transferability from all perspectives for future work.